OpenStudy (anonymous):

Solve |2x-9|+4 less than or equal to7

OpenStudy (anonymous):

this is what I have so far |2x-9|+4<=7 First I subtract 4. |2x-9|<=3 2x+9<=3 Then,subtract 9 both sides. 2x<=-6 2x/2<=-6/2 x<=-3 |2x-9|+4>=7 I subtract 4 from both sides. |2x-9|>=3 |2x-9|>=-3 2x-9>=-3 I add 9 to both sides. 2x>=6

OpenStudy (anonymous):

O.o??

OpenStudy (anonymous):

Does that makes sense?

OpenStudy (anonymous):

What am a little confused if the two answer the negative and the positive are <= <= or <= and >=?

OpenStudy (anonymous):

x<=-3 x>=6 That's the answer. Just like that.

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

oh , I got -6

OpenStudy (anonymous):

O.o?

OpenStudy (anonymous):

(-3)-9=-12 -12/2 =-6

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Why -6? Just x>=6 Sorry, but I don't see your doubt.

OpenStudy (anonymous):

(-3)-9=-12 -12/2 =-6

OpenStudy (anonymous):

my bad. Is x>=3 and x<=-3

OpenStudy (anonymous):

You said: "2x-9>=-3 I add 9 to both sides. 2x>=6" Then we divide by 3 both sides and we get x>=3

OpenStudy (anonymous):

??? ok this is what I have x<-3 and x>-6

OpenStudy (anonymous):

Where does the -6 comes from?

OpenStudy (anonymous):

I made some changes on my calculations may be thats why

OpenStudy (anonymous):

the negative part 2x+9>-3

OpenStudy (anonymous):

2x-9>=-3 2x+9<=3 Those are the equations from which we start. Not 2x+9>-3

OpenStudy (anonymous):

Mmm... ok whhen you have brackets |2x-9| dont you make it positive? I think thats where I got confused

OpenStudy (anonymous):

I started with 2x+9<3 and 2x+9>-3

OpenStudy (anonymous):

It must be: 2x-9<3 and 2x-9>-3

OpenStudy (anonymous):

oh so you dont make whats inside the brackets positive?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

When you change the inequality you have to change the sign only on ONE side

OpenStudy (anonymous):

ok so appliying what you said my answerr is x<6 and x>3

OpenStudy (anonymous):

Excellent!

OpenStudy (anonymous):

Any doubt?

OpenStudy (anonymous):

nope thank you! :)

OpenStudy (anonymous):

Thanks to you for that medal! Glad to help you!

OpenStudy (anonymous):

:)) ohh also I forgot to ask

OpenStudy (anonymous):

how would you describe that on a graph?

OpenStudy (anonymous):

that is positive?? may be

OpenStudy (anonymous):

x<6 and x>3  Like that?